OpenStudy (anonymous):

Question : 7 - Signals: The continuous time Convolution Integrals \(y(t) = cos( \pi t)[u(t+1) - u(t-1)] * u(t)\) is ?? \(*\) denotes Convolution here..

OpenStudy (anonymous):

Why I am getting 0, I don't know..

OpenStudy (anonymous):

You are typing or not?? Or you have been typing since my birth.. :P

OpenStudy (anonymous):

@amistre64 your one look may clear my doubt..

OpenStudy (badhi):

For \(t\leq 1\) the convolution become zero. Otherwise, It does not become zero.

OpenStudy (anonymous):

I want the total answer from -infinity to +infinity..

OpenStudy (amistre64):

convolutions are not in my repetiore ... and neither is spelling repetiore apparently

OpenStudy (anonymous):

Oh, it is okay @amistre64 .. That's why I said "may"

OpenStudy (anonymous):

By distribution property, firstly I have separated the two convolutions..

OpenStudy (badhi):

for \(t\leq -1\) , \(y(t) = 0\) for \(-1<t\leq1\), \(y(t)=\int \limits_{-1}^{t} \cos (\pi \tau)\;d\tau\) for \(1<t\), \(y(t)=\int \limits_{-1}^{1} \cos(\pi \tau)\; d\tau\)

OpenStudy (anonymous):

How you are analyzing it, can you tell me in detail?

OpenStudy (anonymous):

\[y(t) = \cos(\pi t) \cdot u(t+1) * u(t) - \cos( \pi t) \cdot u(t-1) * u(t)\]

OpenStudy (anonymous):

u(t) can be replaced but It will convert the function in the form of integral from -infinity to t..

OpenStudy (badhi):

Sure, take \(g(t)=\cos(\pi t)[u(t+1)-u(t-1)]\) and \[h(t)=u(t)\] take \[y(t)=\int_{-\infty}^\infty g(\tau)h(t-\tau) d\tau\] graph of g(t) is in the picture

OpenStudy (anonymous):

\[y(t) = \int\limits_{- \infty}^t \cos( \pi t) \cdot u(t+1) \cdot dt - \int\limits_{- \infty}^t \cos( \pi t) \cdot u(t-1) \cdot dt \\y(t) = \int\limits_{- 1}^t \cos( \pi t)\cdot dt - \int\limits_{1}^t \cos( \pi t) \cdot dt\]

OpenStudy (anonymous):

And then I do: \[y(t) = \int\limits_{-1}^1 \cos( \pi t) \cdot dt + \int\limits_1^t \cos(\pi t) dt - \int\limits_1^t \cos(\pi t) dt \\y(t) = \int\limits_{-1}^1 \cos( \pi t) \cdot dt\]

OpenStudy (anonymous):

Where am I doing wrong??

OpenStudy (badhi):

You have only cosidered \[\cos(\pi t)[u(t+1)-u(t-1)]\] which is just integration. Where is \(u(t)\)??? The final convolution should be \[\int_{-\infty}^\infty (\cos(\pi \tau)[u(\tau+1)-u(\tau-1)] )\ast u(t-\tau) d\tau\] For this kind of problem a graphical method is more suitable

OpenStudy (anonymous):

u(t) becomes integration..

OpenStudy (anonymous):

See, here if we convolve, u(t) with u(t), then: \[\int\limits_{- \infty} ^{\infty}u(\tau) \cdot u(t - \tau) d \tau \implies \int\limits_{-\infty}^t u(\tau) d(\tau)\]

OpenStudy (anonymous):

This means convolving with u(t) will introduce just an integral from -infinity to t..

OpenStudy (anonymous):

Are you getting what I am try to say??

OpenStudy (badhi):

But there is a \(\cos(\pi t)\) attached to u(t+1) and u(t-1) with it. So it is not just convolving u(t) with u(t)...

OpenStudy (anonymous):

Okay, I will clear it when I will get some time, you explain your method..

OpenStudy (anonymous):

yes, I can see that graph, cos(pi t) multiplied with rectangular function will given that.. :)

OpenStudy (badhi):

You are assuming that \(g(t)\ast u(t) = \int_{-\infty}^tg(\tau) d\tau\) which is wrong

OpenStudy (anonymous):

And u(t) with u(t) was just an example, in general you can take any function.. Why it is wrong??

OpenStudy (anonymous):

Okay, explain me how that is wrong, I am listening..

OpenStudy (anonymous):

@myininaya your eyes will be definitely need here.. (Not physically though).. :)

OpenStudy (anonymous):

*needed..

OpenStudy (anonymous):

@BAdhi why you have stopped??

OpenStudy (anonymous):

Explain that..

OpenStudy (badhi):

Sorry, I was wrong about \(g(t)\ast u(t) = \int_{-\infty}^t g(\tau)d\tau\). It is a correct property. You can use this straight forward to find the answer but do not expand the \(cos(\pi t)[u(t+1)-u(t-1)]\). This denotes a region of cos(pi t) from -1 to 1. so the convolution \(y(t) = \int_{-\infty}^t g(\tau) \; d\tau \) What you have made a mistake is taking \(\tau\) as \(t\). t here is like a constant in the integral. So for example if t<-1, the y(t) integral will be zero since \(g(\tau)\) is from -1 to 1. But it will not be 0 if t>-1. Since it denotes the area under the g(t) which is shown in the picture below

OpenStudy (anonymous):

That seems to be right.. I was thinking the same that I should work with \(\tau\) there.. :)

OpenStudy (anonymous):

I got it now.. :) A real thanks.. :)